OpenStudy (anonymous):

How to find sin 300 without calculator

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

sin 300 = sin(360-60) =-sin60 = -sqrt(3)/2= -0.866(appx)

OpenStudy (anonymous):

sin(360-theta) is - sintheta?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sin(360+-theta) is +- sintheta

OpenStudy (anonymous):

oh depending upon the quadrant theta lies in

ganeshie8 (ganeshie8):

u knw \(\sin(A\pm B)\) formula right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It would be a nightmare suppose apply it to sin 5000

ganeshie8 (ganeshie8):

\(\sin(360\pm \theta) = \sin 360 \cos \theta \pm \cos 360 \sin \theta = 0 \pm \sin \theta \)

OpenStudy (anonymous):

such questions like sin 4000 and find cos 3500 come in exam so this identities won't be useful

OpenStudy (anonymous):

i mean that was just an example, i know it won't be a perfect value

ganeshie8 (ganeshie8):

you may reduce sin(10^10000) easily, use below : \(\large \sin(x \pm 360n) = \sin(x) \)

OpenStudy (anonymous):

for less than 360

OpenStudy (anonymous):

and for other 5 ratios

ganeshie8 (ganeshie8):

yes you can make it less than 360, and after that you can use other identities to compute the numerical value if possible. since sin is a periodic function of period 360 degrees, adding/subtracting 360 degrees will not change its value

ganeshie8 (ganeshie8):

may be lets work an example and see how this works ?

OpenStudy (anonymous):

okay!

ganeshie8 (ganeshie8):

find the value of \(\large \sin(360000000030) = ?\)

OpenStudy (anonymous):

that goes into 360 perfectly

ganeshie8 (ganeshie8):

nope, look carefully again..

ganeshie8 (ganeshie8):

\[\sin(360000000030) = \sin(36000000000 + 30) \\= \sin (360n + 30) = \sin(30) = \dfrac{1}{2}\]

OpenStudy (nincompoop):

sin (9^9^9^9^9)

OpenStudy (anonymous):

oh i see

ganeshie8 (ganeshie8):

that looks like a very interesting problem

ganeshie8 (ganeshie8):

\[\large \sin \left(9^{9^{9^{9^9}}}\right) = \sin \left(9^{9^{9^{9^9}}} \mod 360\right) \]

ganeshie8 (ganeshie8):

we need to find the remainder when 9^9^9^9 is divided by 360

OpenStudy (anonymous):

yes i am trying but without calc how i would do i mean i can but still

ganeshie8 (ganeshie8):

may be, il let you figure it out, the answer is : \(\sin(9)\)

OpenStudy (anonymous):

ok for tan 225

ganeshie8 (ganeshie8):

whats the period of tan ?

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

180 in degrees

ganeshie8 (ganeshie8):

so adding/subtracting 180 degrees doesn't change the value of tan

ganeshie8 (ganeshie8):

it takes u back to ur older value

ganeshie8 (ganeshie8):

so, \(\tan (x \pm 180) = \tan (x)\)

ganeshie8 (ganeshie8):

use it and reduce

OpenStudy (anonymous):

i got it for sin and cos adding and subtracting values from 360 will give the same answer and subtracting it from 180 would give " - " of the answer

ganeshie8 (ganeshie8):

kindof, but its not a good idea to cram our heads with too many formulas

OpenStudy (anonymous):

i will come back after some time sry

ganeshie8 (ganeshie8):

i think remembering below is sufficient : 1) period : \(\sin (x \pm 360) = \sin (x)\) \(\cos (x \pm 360) = \cos (x)\) \(\tan (x \pm 180) = \tan (x)\) 2) angle sum/difference formulas

OpenStudy (nincompoop):

laughing out loud

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i dont get it at all say for sin(x ± 360) = sin (x) .....what do you mean. like if you have 22..you to take away 360 or add. Im confused

ganeshie8 (ganeshie8):

You're correct monika :) take away 360 or add 360, the sin value will stay same.

ganeshie8 (ganeshie8):

for example : sin(30 + 360) = sin(30) sin(30 - 360) = sin(30

OpenStudy (anonymous):

How would you know which one to do..take away or add? or do you just put the sign. i thought it came about a different answer...thank you!

ganeshie8 (ganeshie8):

hmm we won't do anything unless it is required

ganeshie8 (ganeshie8):

to find \(\large \sin(390)\) , you will need to subtract 360

ganeshie8 (ganeshie8):

to find \(\large \sin (-330)\) , you will need to add 360

ganeshie8 (ganeshie8):

depends on what we need to find...

OpenStudy (anonymous):

so sin(390) would end up .. sin( 390 - 360) = sin(30)?

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

you're free to subtract as many 360's as you wish - the value will not change.

OpenStudy (anonymous):

Thank you so so much!

OpenStudy (anonymous):

sin(90+theta) = cos theta cos(90+theta)= - sin theta cos(180+theta)= -cos theta tan(180 + theta) = tan theta